OpenStudy (anonymous):

express the exact value of of csc 60 degrees with a rational denominator. please show all work.

OpenStudy (cwrw238):

csc 60 = 1 / sin 60 = 1 / (sqrt3/2) = 2 / sqrt3

OpenStudy (cwrw238):

can you continue from here - rationalize the denominator

OpenStudy (cwrw238):

multiply top and bottom by sqrt3

OpenStudy (anonymous):

im getting confused csc 60= 1 how? also how comes u must use sin and square root can u explain it a bit more please.

OpenStudy (cwrw238):

no its a trig identity csc x = 1/sin x

OpenStudy (cwrw238):

sin 60 = sqrt3 / 2 so csc 60 = 1 / (sqrt3/2) = 2 / sqrt3

OpenStudy (cwrw238):

rationalizing the denominator 2 / sqrt3 * sqrt3 / sqrt3 = 2sqrt3 / 3

OpenStudy (anonymous):

okay i understood the sin 60 part. can u help me with the rationalizing denominator again. what did u do with the 2?